OpenStudy (anonymous):

The length of a rectangle, l, is six meters more than its width, w. If the perimeter of the rectangle is 36 meters, what is the length of the rectangle, in meters? A. 12 B. 6 C. 15 D. –12

OpenStudy (fifciol):

2l + 2a= 36 a=l-6 2l + 2(l-6)=36 4l-12=36 4l=48 l=12 ans A

OpenStudy (fifciol):



OpenStudy (anonymous):

thank you